# KunquDB: An Attempt for Speaker Verification in the Chinese Opera Scenario

### v1

@@ Abstract
This work aims to promote Chinese opera research in both musical and speech domains, with a primary focus on overcoming the data limitations. We introduce KunquDB, a relatively large-scale, well-annotated audio-visual dataset comprising $339$ speakers and $128$ hours of content. Originating from the Kunqu Opera Art Canon (Kunqu yishu dadian), KunquDB is meticulously structured by dialogue lines, providing explicit annotations including character names, speaker names, gender information, vocal manner classifications, and accompanied by preliminary text transcriptions. KunquDB provides a versatile foundation for role-centric acoustic studies and advancements in speech-related research, including Automatic Speaker Verification (ASV). Beyond enriching opera research, this dataset bridges the gap between artistic expression and technological innovation. Pioneering the exploration of ASV in Chinese opera, we construct four test trials considering two distinct vocal manners in opera voices: stage speech (ST) and singing (S). Implementing domain adaptation methods effectively mitigates domain mismatches induced by these vocal manner variations while there is still room for further improvement as a benchmark.
@@ 1 Introduction
The scarcity of detailed annotations poses a significant obstacle for numerous research tasks on opera data. This obstacle is particularly pronounced for tasks requiring comprehensive annotations, including automatic speaker recognition for speaker label prediction, Automatic Speech Recognition (ASR) for text transcription retrieval, speaker diarization for role detection, as well as speech and singing voice synthesis. Meanwhile, speech-related research predominantly focuses on conventional speech. Existing open-source models designed for various speech tasks, such as ASR and speaker diarization, exhibit inadequate robustness when applied to opera data. The complex acoustic characteristics in opera voices provide a diverse testing ground for evaluating the robustness of speech models. The absence of automated tools further obstructs large-scale data collection and cleaning, restricting access to diverse and abundant datasets. This dilemma creates a cycle that impedes progress in data availability, hindering the development of advanced tools for digitizing opera research.
@@
In response to the challenges posed by insufficient data and limitations of existing models in the field of Chinese opera, our primary objective is to create a symbiotic relationship between data and models. To achieve this, we present a comprehensive and publicly accessible audio-visual dataset characterized by its richness and scale. This resource is designed to lay the groundwork for developing specialized automated tools applicable to Chinese opera, thereby facilitating advancements in the study of this art form. Narrowing down from the landscape of Chinese traditional opera, we focus on the exquisite domain of Kunqu Opera. Reputed as the mother of Chinese operas, Kunqu Opera boasts a history spanning over 600 years [10], giving rise to numerous operas, including Beijing Opera. In alignment with [5], we selectively choose classic and authoritative audio-visual materials sourced from the Kunqu Opera Art Canon (Kunqu yishu dadian) [^1] Note: After purchasing the book, we negotiated with the publisher and secured their authorization for its utilization in Kunqu Opera research. The publisher explicitly stated that the book’s digital resource can be employed solely for scholarly or research endeavors upon the approval of the publisher. It may not be illegally disseminated or used for commercial purposes. [39] to ensure both quantity and quality. The source video undergoes sentence-level segmentation, generating preliminary text transcriptions. Subsequently, we proceed with speaker annotation and explicitly categorize each utterance as either stage speech (ST) or singing (S) based on vocal manner. Ultimately, KunquDB [^2] https://hualizhou167.github.io/KunquDB, the curated audio-visual dataset comprises 339 performers, totals approximately 128 hours, with stage speech and singing voices each constituting about half of the dataset. As an audio-visual dataset, it is applicable in various scenarios, including ASV, ASR, speaker diarization, singing voice synthesis, person re-identification and multi-modal understanding.
@@
  • We curate KunquDB, a comprehensive audio-visual dataset specifically tailored for Kunqu Opera. Its large scale effectively mitigates data shortages and fosters a positive feedback loop between data and models.
@@ 3.1 Chinese Opera
Due to the lack of automatic tools for data collection and cleaning, existing opera datasets [5, 16, 12, 6] are limited in scale and annotation richness. Typically spanning only a few hours [16, 6] and offering plain labels [6], they are insufficient for downstream recognition tasks like ASV and singing speech recognition in the field of Chinese opera. In the intersection of opera and speech-related research, most efforts are focused on synthesis, with reliance on the only publicly accessible, well-annotated yet small-scale dataset, "Jingju a cappella singing" [12]". It serves as the basis for subsequent neural network-based opera synthesis by [41, 42, 49, 25]. While [41] and [42] pioneer neural network-based synthesis using the DurIAN [46] framework, [49] introduces OperaSinger, based on the FastSpeech2 [30] framework, exploring novel data augmentations within this small-scale dataset [12]. In a related vein, [25] attempts to transfer popular singers’ timbre to Chinese opera using the VITS [18] model with the same dataset [12].
@@ 3.2 Automatic Speaker Verification
The neural network backbone, acting as the encoder, extracts frame-level features from the input utterance. This backbone has evolved from structures like 2D Convolutional Neural Networks (CNNs) [37], Time-Delay Neural Networks (TDNNs) [9], and Transformers [13]. Currently, prevalent models are based on 2D CNNs with ResNet [14]. The pooling layer aggregates frame-level features into a fixed-length, utterance-level representation, which is then projected linearly to generate the speaker embedding. Common temporal aggregation techniques include average pooling [38], statistical pooling [33] and attentive pooling [24]. The loss function is the optimized objective during training, such as the Additive Margin Softmax (AM-Softmax) [36] and ArcFace [8]. The scoring strategy, or the back-end model, measures the similarity between enrollment and test utterance embeddings for verification. Typically, cosine similarity or Probabilistic Linear Discriminant Analysis (PLDA) [26] are utilized.
@@ 4.1.2 Why is Manual Labeling Required?
Due to the nature of Kun Opera performances, where the entire stage is often captured rather than close-ups of characters, human faces occupy limited space in the frame. Moreover, performers’ heavy makeup and theatrical costumes further obscure facial features, particularly the waist-length beards (rankou[40] worn by male characters typically completely conceal their mouths. Consequently, conventional pipelines used in  [23, 22], involving face detection, tracking, audio-video synchronization for mouth movement and speech, and face verification, are unsuitable for these opera videos.
@@ 4.1.3 How to Get KunquDB?
The book [39] purchase grants access to the digital source video data in a supplementary disc. It is the user’s responsibility to get the approval from the publisher to conduct research for non-commercial purposes. We provide annotated data, including segment start and end timestamps, along with associated information, including character names, speaker names, and preliminary text transcriptions. The open-source annotations and processing scripts can be accessed and downloaded online\@footnotemark.
@@ 4.2.1 Step 1: Video Segmentation
We utilize VideoSubFinder [^3] https://sourceforge.net/projects/videosubfinder, in conjunction with PaddleOCR [^4] https://github.com/PaddlePaddle/PaddleOCR to extract hardcoded subtitles from source videos, yielding start and end timestamps for each dialogue line and corresponding text transcriptions. Using ffmpeg [^5] https://ffmpeg.org, we then segment the videos into clips based on the acquired timestamps, resulting in individual video clips for each dialogue line.
@@ 4.3 Dataset Statistics
Table 1 summarizes key statistics for the KunquDB dataset, differentiating stage speech (ST) and singing (S) categories. The dataset comprises $60,066$ ST utterances and $17,902$ S utterances, contributed by $288$ speakers for both ST and S data, $50$ exclusively providing ST data, and $1$ exclusively offering S data. Additionally, there are $339$ videos featuring both ST and S, along with $5$ exclusively for ST and $2$ for S. Fig. 2 visually represents the distribution of utterance lengths and speakers enacting role types.
@@ 4.5 Trial Construction
  • Undifferentiated Trial: No distinction between enrollment and test utterance types regarding vocal manner; samples are randomly chosen from either stage speech or singing.
@@ 5.1 Domain Discrepancy Adversarial Learning
We leverage an attention mechanism to disentangle domain-related features, expressed as $\mathbf{f_{domain}}$, induced by different vocal manners from the extracted features $\mathbf{f}$ by the backbone model. Next, we refine speaker-specific features, $\mathbf{f_{id}}$, by filtering out $\mathbf{f_{domain}}$. Following this, both $\mathbf{f_{domain}}$ and $\mathbf{f_{id}}$ undergo pooling and fully connected layers, producing the domain embedding $\mathbf{z_{domain}}$ for domain classification and speaker ID embedding $\mathbf{z_{id}}$ for speaker classification. Further, we employ a gradient reversal layer (GRL) before an auxiliary domain classifier to eliminate domain influence from $\mathbf{z_{id}}$ through adversarial learning.
@@ 6.1.1 Dataset
We pretrain the model on VoxBlink2 with over 16,000 hours of audio data from 110k speakers. Thereupon, we fine-tune the model using KunquDB’s training set. Evaluation is performed on the KunquDB test set.
@@ 6.1.5 Experimental Results
    Model M0 shows weak robustness on Kunqu data, performing best in the ST-domain due to its exclusive pretraining on speech data. However, its performance is still markedly inferior to its excellent performance on regular speech test sets, often below $1\%$ EER.
@@
    Models generally perform best when enrollment and testing vocal manner match, either in the S or ST domain. Nevertheless, their performance notably decreases in cross-domain scenarios, indicating the difficulty in extracting domain-agnostic speaker embeddings.

### v2

@@ Abstract
This work aims to promote Chinese opera research in both musical and speech domains, with a primary focus on overcoming the data limitations. We introduce KunquDB, https://hualizhou167.github.io/KunquDB, a relatively large-scale, well-annotated audio-visual dataset comprising $339$ speakers and $128$ hours of content. Originating from the Kunqu Opera Art Canon (Kunqu yishu dadian), KunquDB is meticulously structured by dialogue lines, providing explicit annotations including character names, speaker names, gender information, vocal manner classifications, and accompanied by preliminary text transcriptions. KunquDB provides a versatile foundation for role-centric acoustic studies and advancements in speech-related research, including Automatic Speaker Verification (ASV). Beyond enriching opera research, this dataset bridges the gap between artistic expression and technological innovation. Pioneering the exploration of ASV in Chinese opera, we construct four test trials considering two distinct vocal manners in opera voices: stage speech (ST) and singing (S). Implementing domain adaptation methods effectively mitigates domain mismatches induced by these vocal manner variations while there is still room for further improvement as a benchmark.
@@ 1 Introduction
The scarcity of detailed annotations poses a significant obstacle for numerous research tasks on opera data. This obstacle is particularly pronounced for tasks requiring comprehensive annotations, including automatic speaker recognition for speaker label prediction, Automatic Speech Recognition (ASR) for text transcription retrieval, speaker diarization for role detection, as well as speech and singing voice synthesis. Meanwhile, speech-related research predominantly focuses on conventional speech. Existing open-source models designed for various speech tasks, such as speaker diarization, exhibit inadequate robustness when applied to opera data. The complex acoustic characteristics in opera voices provide a diverse testing ground for evaluating the robustness of speech models. The absence of automated tools further obstructs large-scale data collection and cleaning, restricting access to diverse and abundant datasets. This dilemma creates a cycle that impedes progress in data availability, hindering the development of advanced tools for digitizing opera research.
@@
In response to the challenges posed by insufficient data and limitations of existing models in the field of Chinese opera, our primary objective is to create a symbiotic relationship between data and models. To achieve this, we present a comprehensive and publicly accessible audio-visual dataset characterized by its richness and scale. This resource is designed to lay the groundwork for developing specialized automated tools applicable to Chinese opera, thereby facilitating advancements in the study of this art form. Narrowing down from the landscape of Chinese traditional opera, we focus on one exquisite domain, Kunqu Opera. Reputed as the mother of Chinese operas, Kunqu Opera boasts a history spanning over 600 years [10], giving rise to numerous operas, including Beijing Opera. In alignment with [5], we selectively choose classic and authoritative audio-visual materials sourced from the Kunqu Opera Art Canon (Kunqu yishu dadian) [^1] Note: After purchasing the book, we negotiated with the publisher and secured their authorization for its utilization in Kunqu Opera research. The publisher explicitly stated that the book’s digital resource can be employed solely for scholarly or research endeavors upon the approval of the publisher. It may not be illegally disseminated or used for commercial purposes. [39] to ensure both quantity and quality. The source video undergoes sentence-level segmentation, generating preliminary text transcriptions. Subsequently, we proceed with speaker annotation and explicitly categorize each utterance as either stage speech (ST) or singing (S) based on vocal manner. Ultimately, KunquDB [^2] https://hualizhou167.github.io/KunquDB, the curated audio-visual dataset comprises 339 performers, totals approximately 128 hours, with stage speech and singing voices each constituting about half of the dataset. As an audio-visual dataset, it is applicable in various scenarios, including ASV, ASR, speaker diarization, singing voice synthesis, person re-identification and multi-modal understanding.
@@
  • We curate KunquDB\@footnotemark, a comprehensive audio-visual dataset specifically tailored for Kunqu Opera. Its large scale effectively mitigates data shortages and fosters a positive feedback loop between data and tool models.
@@ 3.1 Chinese Opera
Due to the lack of automatic tools for data collection and cleaning, existing opera datasets [5, 16, 12, 6] are limited in scale and annotation richness. Typically spanning only a few hours [16, 6] and offering plain labels [6], they are insufficient for downstream recognition tasks like ASV and singing speech recognition in the field of Chinese opera. In the intersection of opera and speech-related research, most efforts are focused on synthesis, with reliance on the only publicly accessible, well-annotated yet small-scale dataset, "Jingju a cappella singing" [12]. It serves as the basis for subsequent neural network-based opera synthesis by [41, 42, 49, 25]. While [41] and [42] pioneer neural network-based synthesis using the DurIAN [46] framework, [49] introduces OperaSinger, based on the FastSpeech2 [30] framework, exploring novel data augmentations within this small-scale dataset [12]. In a related vein, [25] attempts to transfer popular singers’ timbre to Chinese opera using the VITS [18] model with the same dataset [12].
@@ 3.2 Automatic Speaker Verification
The neural network backbone, as the encoder, extracts frame-level features from the input utterance. This backbone has evolved from architectures like 2D Convolutional Neural Networks (CNNs) [37], Time Delay Neural Networks (TDNNs) [9], and Transformers [13]. Currently, 2D CNNs with ResNet [14] are the most widely adopted. The pooling layer aggregates frame-level features into a fixed-length, utterance-level representation, which is then projected linearly to generate the speaker embedding. Common temporal aggregation techniques include average pooling [38], statistical pooling [33] and attentive pooling [24]. The loss function is the optimized objective during training, such as the Additive Margin Softmax (AM-Softmax) [36] and ArcFace [8]. The scoring strategy, or the back-end model, measures the similarity between enrollment and test utterance embeddings for verification. Typically, cosine similarity or Probabilistic Linear Discriminant Analysis (PLDA) [26] are utilized.
@@ 4.1.2 Why is Manual Labeling Required?
Due to the nature of Kun Opera performances, where the entire stage is often captured rather than close-ups of characters, human faces occupy limited space in the frame. Moreover, performers’ heavy makeup and theatrical costumes further obscure facial features, particularly the waist-length beards (rankou[40] worn by male characters typically completely conceal their mouths. Consequently, conventional pipelines, as used in  [23, 21], involving face detection, tracking, verification, and audio-video synchronization for mouth movement and speech, are unsuitable for these opera videos.
@@ 4.1.3 How to Get KunquDB?
The book [39] purchase grants access to the digital source video data in a supplementary disc. It is the user’s responsibility to get the approval from the publisher to conduct research for non-commercial purposes. We provide annotated data, including segment start and end timestamps, along with associated information, such as character names, speaker names, and preliminary text transcriptions. The open-source annotations and processing scripts can be accessed and downloaded online\@footnotemark.
@@ 4.2.1 Step 1: Video Segmentation
We utilize VideoSubFinder [^3] https://sourceforge.net/projects/videosubfinder, in conjunction with PaddleOCR [^4] https://github.com/PaddlePaddle/PaddleOCR to extract hardcoded subtitles from source videos, yielding timestamps for each dialogue line and corresponding text transcriptions. Using ffmpeg [^5] https://ffmpeg.org, we then segment the videos into clips based on the acquired timestamps, resulting in individual video clips for each dialogue line.
@@ 4.3 Dataset Statistics
Table 1 summarizes key statistics for the KunquDB dataset, differentiating stage speech (ST) and singing (S) categories. The dataset contains $60,066$ ST utterances and $17,902$ S utterances, contributed by $288$ speakers for both ST and S data, $50$ exclusively providing ST data, and $1$ exclusively offering S data. Additionally, there are $339$ videos featuring both ST and S, along with $5$ exclusively for ST and $2$ for S. Fig. 2 visually represents the distribution of utterance lengths and speakers enacting role types.
@@ 4.5 Trial Construction
  • Undifferentiated Trial: No distinction between enrollment and test utterance regarding vocal manner categories; samples are randomly chosen from either stage speech or singing.
@@ 5.1 Domain Discrepancy Adversarial Learning
We leverage an attention mechanism to disentangle domain-related features $\mathbf{f_{domain}}$ induced by different vocal manners from the feature map $\mathbf{f}$ extracted by the backbone model. Next, we refine speaker-specific features, $\mathbf{f_{id}}$, by filtering out $\mathbf{f_{domain}}$. Following this, both $\mathbf{f_{domain}}$ and $\mathbf{f_{id}}$ undergo pooling and fully connected layers, producing the domain embedding $\mathbf{z_{domain}}$ for domain classification and speaker ID embedding $\mathbf{z_{id}}$ for speaker classification. Further, we employ a gradient reversal layer (GRL) before an auxiliary domain classifier to eliminate domain influence from $\mathbf{z_{id}}$ through adversarial learning.
@@ 6.1.1 Dataset
We pretrain the model on VoxBlink2 [21] with over 16,000 hours of audio data from 110k speakers. Thereupon, we fine-tune the model using KunquDB’s training set. Evaluation is performed on the KunquDB test set.
@@ 6.1.5 Experimental Results
    Model M0 shows weak robustness on Kunqu data, performing best in the ST-domain due to its exclusive pretraining on speech data. Nevertheless, its performance is still markedly inferior to its excellent performance on regular speech test sets, often below $1\%$ EER.
@@
    Models generally perform best when enrollment and test utterances share the same vocal manner, whether in the S or ST category. However, their performance notably declines in cross-domain scenarios, indicating the difficulty in extracting domain-agnostic speaker embeddings.